OpenStudy (anonymous):

Consider 3 trials, each having the same probability of success. Let X denote the total number of successes in these trials. If E[X]=1.8 what is the largest possible value of P{X=3}?

OpenStudy (anonymous):

The answer is .6 and 0

OpenStudy (anonymous):

Ok?

OpenStudy (anonymous):

Use Chebishohv. inequality

OpenStudy (anonymous):

P(abs(3-1.8)>=ksigma)<=1/k2

OpenStudy (anonymous):

(1)P(1) + (2)P(2) + (3)P(3) = 1.8 If we let P(2) = 0, then (1)[1 - P(3)] + (3)P(3) = 1.8 will give P(3) = 0.4 which is the highest allowable for P(3)

OpenStudy (anonymous):

Let me check P(1)=0 2(1-p)+3p=1.8 p<0 so P(2)=0 gives Maximum P. Great